OpenStudy (anonymous):

Can someone help me with this??? Which of the following is an equation of the line that passes through the point (1, 1) and has a slope of 2? A. y = 2x – 1 B. y = 2x + 1 C. y = –2x – 1 D. y = –2x + 1

OpenStudy (anonymous):

Recall y =mx+b m = slope x is a number in a domain. and b is the y intercept. It says that when x =1 then y =1 so you can write 1=2(1)+b then solve for b and write out what y=mx+b.

OpenStudy (anonymous):

Ok I got -2=b. But it doesn't answer the question because it's not one of the choices.

OpenStudy (anonymous):

1-2=b b=-1

OpenStudy (anonymous):

Yes, recheck your math again. You should get -1.